29
Feb
2012
16:05

We just suffered a problem with one of our core routers in MHA, resulting in services in that datacentre being off for 11 minutes. Investigations as to the root cause are under way.

Posted by Kevin Hones

17
Feb
2012
12:07

All BT based ADSL connections briefly glitched just now. Apologies for this – we are awaiting an explanation from our wholesale supplier.

Posted by Kevin Hones

10
Feb
2012
10:09

All Be connections are now running normaly.

Posted by Phil McGillivray

10
Feb
2012
08:15

It appears engineers have finished restoring configurations. Most customers are now back on line a further update will be issued at 10:00. We apologise if this outage has caused you problems.

Posted by Phil McGillivray

10
Feb
2012
08:06

Be lines have suffered a major outage. Service is being restored and a further update will be available in 15 Minutes. This problem was caused by overnight changes made to Be core routers.

Posted by Phil McGillivray

Current Status

    GoodNo current issues

Status Archive